What is the total amount of energy needed (in unit of J) to ionize one electron from the outermost subshell of 10 billion atoms of X in a photoelectron spectroscopy experiment?
Answer #1

ANSWER:-

Photoelectron spectroscopy (PES) is the energy measurements of photoelectrons emitted from solids, gases, or liquids by the photoelectric effect. Depending on the source of ionization energy, PES can be divided accordingly into Ultraviolet Photoelectron Spectroscopy (UPS) and X-ray Photoelectron Spectroscopy (XPS). The source of radiation for UPS is a noble gas discharge lamp, usually a He discharge lamp. For XPS, also referred to as Electron Spectroscopy for Chemical Analysis (ESCA), the source is high energy X-rays (1000-2000 eV). Furthermore, depending on the source of the ionization energy, PES can probe either valence or core electrons. UPS, which uses the energy of ultraviolet rays (<41 eV)(10-45eV) will only be sufficient to eject electrons from the valence orbitals, while the high energy X-rays used in XPS can eject electrons from the core and atomic orbitals

that is when XPS is used then energy required = 1000-2000eV that is 2000eV= 1.6x 10-19 Joule that is 1.6x2000x10-19 = 3.2 x 10-16 joule

or when UPS is used than 45eV = 1.6x 10-19 x 45 = 7.2 x 10-18 joule.